# Exchange control in a MOS double quantum dot made using a 300 mm wafer process

### v2

@@ Abstract
From the $T_{2}^{*}$ data shown in Fig. 3(f), we observe a rapid increase in coherence as the detuning increases from zero, indicative of a low $\delta\varepsilon_{\text{rms}}$. The extracted value of $\delta\varepsilon_{rms}=5.4\pm 0.1~{}\upmu$eV, obtained over a measurement time of 79 h, is at the state-of-the-art [21, 33, 18, 47, 48, 43], and can be attributed to the low charge noise achieved for samples using this 300 mm process [10]. As the detuning increases further, where $J<\Delta E_{z}$, we observe that noise in $\Delta E_{z}$ dominates (due to 29Si nuclear spins), leading to a relatively constant $T_{2}^{*}$. From this saturation value of $T_{2}^{*}=0.28\pm 0.04~{}\upmu$s, we extract $\sigma_{\text{HF}}=\sqrt{2}\pi\delta\Delta E_{\text{z,rms}}/(g\mu_{B})=0.13\pm 0.02$ mT. Note that we assume the Zeeman energy fluctuations are dominated by the Overhauser field rather than noise in the g factor difference.
The entangling two-qubit gate achieved between the spin qubits under the exchange interaction depends on the ratio $J/\Delta E_{z}$, tending to a $\sqrt{\text{SWAP}}$ operation as $J\gg\Delta E_{z}$ or a C-PHASE when $J\ll\Delta E_{z}$, though any gate within this set parameterised by $J/\Delta E_{z}$ can be used as the building block for a quantum error correcting code such as the surface code [49]. Defining the gate quality factor as $Q_{\rm gate}=T_{2}^{*}\Omega$ (i.e. the number of periods before the amplitude of oscillations decays by $1/e$), we find $Q_{\rm gate}\gtrapprox 10$ in the region $J/\Delta E_{z}=2.1-3.2$. This provides an upper bound estimate on the achievable two-qubit gate fidelity using the approximation $\mathcal{F}\approx 1-1/4Q_{\rm gate}\lessapprox 98\%$ [50]. To implement error-correctable two-qubit gates this fidelity would need to surpass $99\%$ [1, 2, 3], which could be achieved using isotopically enriched silicon. In the next section, we extend the coherence time using spin refocusing techniques.
Figure 4: |||| Echo sequence. a, Bloch sphere representation of the odd-parity two spin sub-space, indicating the rotation axes, $\hat{Z^{\prime}}$ and $\hat{X^{\prime}}$ and their angular deviation from the nominal $\hat{Z}$ axis defined by |$\ket{S}$ and |$\ket{T_{0}}$. b Schematic of the exchange echo sequence. c Echo signal as a function of free evolution time difference $\tau_{2}-\tau_{1}$. d Echo amplitude as a function of total free evolution time $\tau_{2}=\tau_{1}$. Data (purple dots) and exponential decay fit (orange line).
@@
In the exchange echo sequence, after initialising a |$\ket{\text{S}(1,1)}$ state and applying a $X^{\prime}_{\pi/2}$ rotation, the two-electron system dephases under the effect of charge noise for a time $\tau_{1}$. The free evolution occurs at a detuning point where $J/\Delta E_{z}=12.8\pm 1.6$ where we measured $T_{2}^{*}=43\pm 3~{}$ns (see Fig 3(f)). We then refocus the spins by applying a $X^{\prime}_{\pi}$ rotation and let the system evolve for $\tau_{2}$ until a second $X^{\prime}_{\pi/2}$ rotation maps the resulting state to the |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ axis. We extract the amplitude of the echo by fitting the signal in the $\tau_{2}-\tau_{1}$ domain (see Fig. 4(c)), and plot its value as a function of total free evolution time $\tau_{1}+\tau_{2}$. Fig. 4(d) shows the echo amplitude decays exponentially with total time ($\tau_{2}+\tau_{1}$), yielding a characteristic $T_{2,J}^{\text{echo}}=0.42\pm 0.02~{}\upmu$s which corresponds to an order of magnitude increase in the coherence time. From our fit to Eq. (3), we extract a magnetic-noise-limited $T_{2,\Delta E_{z}}^{*}=3.3~{}\upmu$s at this set point, indicating that sources other than magnetic noise limit $T_{2,J}^{\text{echo}}$. We speculate that residual high frequency charge noise limits the echo coherence [38], and may include factors such as the effect of the rf tone which is on throughout the control sequence.
@@ Discussion
We have demonstrated exchange control, which forms the basis for two-qubit gates between spin qubits, in an MOS DQD device fabricated using a 300 mm wafer process on natural silicon. The quality of the manufacturing manifests in low detuning noise and relatively long $T_{2}^{*}$ for natural silicon in a MOS device. These results highlight the benefits of industrial manufacturing and encourage follow-up studies in isotopically enriched Si samples with larger numbers of QDs. Introducing dedicated gates to control primarily the exchange strength over a wider range could be used to (i) enable symmetric exchange pulses and (ii) reduce $J/\Delta E_{z}$ well below 1, which should lead to an overall reduction in sensitivity to charge noise. Furthermore, the new dispersive readout method demonstrated here using an rf-cascade, expands the portfolio of rf readout methods and provides a solution to the relatively low sensitivity of $in-situ$ dispersive sensing in planar MOS devices. These results may stimulate experiments in which the cascade effect is propagated and readout at a distance which could be achieved by introducing intermediary cascaded QDs. Such an approach would enable enhanced-signal dispersive readout over long distances within dense QD arrays.
@@ Methods
Measurement set-up. We perform the measurements at the base temperature of a dilution refrigerator ($T\sim 10$ mK). We send low-frequency signals through cryogenic low-pass filters with a cut-off frequency of 65 kHz, while we apply pulsed signals through attenuated coaxial lines. Both signals are combined through bias-Ts at the sample PCB (printed circuit-board) level. The PCB was made from RO4003C 0.8 mm thick with an immersion silver finish. For readout, we use radio-frequency reflectometry applied on the ohmic contact of the device. We send radio-frequency signals through attenuated coaxial lines to an on-PCB $LC$ resonator with parallel topology formed by a a coupling capacitor ($C_{\text{c}}$), a 100 nm-thick NbTi superconducting spiral inductor ($L$) and the parasitic capacitance to ground ($C_{\text{p}}$), see Fig. 1(b). We drive the resonator at 512.25 MHz which is the frequency of the system when $G_{\text{S}}$ is well above threshold. The reflected rf signal is then amplified at 4 K and room temperature, followed by quadrature demodulation, from which the amplitude and phase of the reflected signal were obtained (homodyne detection).
Data Availability
Acknowledgements
We acknowledge helpful conversations with J.M. Williams, H. Jnane, A. Seigel, S.C. Benjamin, A.J. Fisher and G. Burkard at Quantum Motion. We also acknowledge technical support from G. Antilen Jacob at the London Centre for Nanotechnology. This work received support from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 951852 (Quantum Large Scale Integration in Silicon); from the Engineering and Physical Sciences Research Council (EPSRC) under grant Nos. (EP/S021582/1), (EP/L015978/1), (EP/T001062/1) and (EP/L015242/1); and from Innovate UK under grant Nos. (43942) and (10015036). T.M. acknowledges support from the Winton Programme for the Physics of Sustainability. M.F.G.Z. acknowledges support from the UKRI Future Leaders Fellowship (MR/V023284/1).
Author contributions
J.F.C.W. conducted the experiments and analysed the results presented in this work with input from R.C.C.L., M.A.F., M.F.G.Z. and J.J.L.M.; J.F.C.W., M.A.F. and R.C.C.L. conducted preliminary experiments; T.M. and G.A.O developed the quantum dot rf simulator in appendix D; T. M. performed and wrote about the simulations under the supervision of D. F. W. and M. F. G. Z.; S.M.P. designed the device under the supervision of M.A.F., M.F.G.Z. and J.J.L.M.; J.J. and S.K. fabricated the device under the supervision of B.G.; F.E.v.H. characterised the superconducting resonator under the supervision of M.F.G.Z.; N.J. maintained the experimental setup; J.F.C.W., M.F.G.Z. and J.J.L.M. wrote the manuscript with input from R.C.C.L., M.A.F., N.J. and S.M.P.; M.F.G.Z. and J.J.L.M. conceived and oversaw the experiment.
Competing Interests
J.F.C.W., R.C.C.L., M.A.F. and M.F.G.Z are inventors on a patent related to this work.
@@ Acknowledgements
APPENDIX A: ELECTROSTATIC REQUIREMENTS FOR RF-DRIVEN ELECTRON CASCADE READOUT
The conditions for rf-driven electron cascade are similar to those stated in a previous report on electron cascade with a proximal charge sensor [30]. It can be understood by considering the electrochemical potentials $\mu_{i}$ for each dot Qi, as well as the given charge configuration $(\text{N}_{\text{Q}_{1}},\text{N}_{\text{Q}_{2}},\text{N}_{\text{Q}_{\text{ME}}})$ where $\text{N}_{\text{Q}_{i}}$ refers to the number of charges in dot $\text{Q}_{i}$, as shown in Fig. 1. For cascade to occur the DQD must be tuned to the $\text{Q}_{1}-\text{Q}_{2}$ inter-dot charge transition, such that
@@
Tuning the quantum dot array into rf-driven electron cascade requires precise control of the electrochemical potentials of the QDs, to the order of 10s of microvolts in the device presented here. The tuning procedure is most clearly demonstrated in the $V_{\text{Q}_{1}}-V_{\text{Q}_{2}}$ gate-voltage-space, as shown in Fig. 5 where in each panel $V_{\text{Q}_{\text{ME}}}$ is varied, bringing the system into and out of cascade. In this configuration, the range of $V_{\text{G}_{\text{S}}}$ bias voltages satisfying the conditions for cascade are given by $V_{\text{G}_{\text{S}}}^{casc}=728.6\pm 0.1~{}$mV, a relatively narrow range of voltages as compared to the 141414~{}14mV addition voltage of $\text{Q}_{\text{ME}}$.
Figure 5: |||| Radio-frequency driven electron cascade. a-e, Charge stability diagrams as a function of the QD gate voltages $V_{\text{Gi}}$ for different $G_{\text{S}}$ values refereed as $V_{\text{G}_{\text{S}}}^{\text{casc}}$. Panel c is the bias utilized in the main part of the manuscript. f-j, Simulated charge stability diagrams using the methods in App. E. k-o Schematic depiction of the electrochemical levels for the corresponding bias conditions, arranged column-wise.
@@
In panels (f-j), we present matching radio-frequency simulations of the triple QD system, see App. APPENDIX D: CHARGE STABILITY DIAGRAM SIMULATIOND, that highlight the enhanced intensity of the interdot charge transition. Further, we supplement the explanation with schematics of the electrochemical levels in each $V_{\text{G}_{\text{S}}}$ conditions in panels k-o.
APPENDIX B: AMPLIFICATION FACTOR
Here, we obtain the expression for the signal amplification factor generated by the cascade process, i.e. Eq. 1 in the main text. To determine the amplification factor, we consider two different charge movement events as seen from the electrode connected to the resonator, in this case, the electron reservoir (R):
@@
To obtain the amplification factor, we first consider the expressions of the signal-to-noise ratio (SNR) for case (1) [24]. Particularly, we consider the system is in the low frequency limit $hf_{\text{rf}}\ll\Delta_{\text{c}}$ (where $f_{\text{rf}}$ is the frequency of the resonator and $\Delta_{\text{c}}=2t_{c}$), the small signal regime $Q_{\text{L}}\Delta C_{\text{Q}}/(2C_{\text{tot}})\ll 1$ (where $Q_{\text{L}}$ is the loaded quality factor of the resonator, $\Delta C_{\text{Q}}$ is the change in quantum capacitance of the system and $C_{\text{tot}}$ the total capacitance of the system) and the large excitation regime, $\alpha_{21}eV_{\text{dev}}\gg\Delta_{\text{c}}$ to ensure an electron tunnels every half cycle of the rf excitation. Here $\alpha_{21}=\alpha_{\text{r,2}}-\alpha_{\text{r,1}}$ is the interdot lever arm as seen from the reservoir and $V_{\text{dev}}$ is the amplitude of the oscillatory voltage arriving at the reservoir. In this case, the SNR is
@@
Here, the resonator is coupled the the $Q_{\text{ME}}$ via a tunnel barrier, rather than a capacitively coupled gate, and hence the induced charge is $(1-\alpha_{\text{r,ME}})e$.
@@
With these facts, we arrive to the expression of the amplification factor by considering the two scenarios. In scenario (1), the AC current generated by charge transfer between Q1 and Q2 is,
@@
In scenario (2), besides the current produced by process (1), we have the current produced by the cascade process adding to a total of,
@@
We therefore arrive to the expression of the amplification factor $A$,
@@
APPENDIX C: MAGNETOSPECTROSCOPY SIMULATION
Figure 6: |||| Magnetospectroscopy simulations. Eigenergies as a fucntion of detuning for $B=0.5$ T (a) and $B=0$ T (b). Singlets in orange triplets in purple. (c) Normalised quantum capacitance versus detuning and magnetic field showing the locatuon of the singlet-triplet minus crossing (dashed white lines).
Here, we describe the simulations of the magnetospectroscopy map in the inset of Fig. 1(c) that allow us to estimate both the tunnel coupling $t_{\text{c}}$ and the electron temperature $T_{\text{e}}$. We utilize the simplified Hamiltonian:
@@
Here, $T_{\text{e}}$ is the DQD temperature and $Z$ is the partition function over all states [51]. We plot the results of the simulations in Fig. 6 where, in panel (a) and (b), we show the energy spectrum of the system as a function of detuning for $B=0.5$ T and 0 T, respectively, and in panel (c), we plot the normalised quantum capacitance of the systems as a function of $\varepsilon$ and $B$. We find that the best match between the data and simulations occurs when $t_{\text{c}}=2.4$ GHz and $T_{\text{e}}=50$ mK.
APPENDIX D: CHARGE STABILITY DIAGRAM SIMULATION
@@
$\mathbf{C_{cc}}=\begin{pmatrix}24.36&-3.39&-1.31\\ -3.39&19.4&-0.279\\ -1.31&-0.279&29.2\\ \end{pmatrix},$
@@
$\mathbf{C_{cv}}=\begin{pmatrix}5.61&0.912&0.00\\ 3.356&2.83&0.00\\ 0.395&0.124&12.6\\ \end{pmatrix}.$
where each element is given in aF. The number of charges on each dot are given by $\vec{N}$, and $\vec{V_{G}}$ denotes the applied gate voltages. Throughout the simulations, we use natural units and set the charge of the electron $e=1$.

### v3

@@ Abstract
From the $T_{2}^{*}$ data shown in Fig. 4(f), we observe a rapid increase in coherence as the detuning increases from zero, indicative of a low $\delta\varepsilon_{\text{rms}}$. The extracted value of $\delta\varepsilon_{rms}=5.4\pm 0.1~{}\upmu$eV, obtained over a measurement time of 0.3 h per trace, is at the state-of-the-art for MOS devices [26, 25, 57], and can be attributed to the low charge noise achieved for samples using this 300 mm process [36], comparable to reports on SiGe heterostructures [61, 62] (see Methods). As the detuning increases further, where $J<\Delta E_{z}$, we observe that noise in $\Delta E_{z}$ dominates (due to 29Si nuclear spins), leading to a relatively constant $T_{2}^{*}$. From this saturation value of $T_{2}^{*}=0.28\pm 0.04~{}\upmu$s, we extract $\sigma_{\text{HF}}=\sqrt{2}\pi\delta\Delta E_{\text{z,rms}}/(g\mu_{B})=30\pm 4~{}\upmu{\rm T}$. Note that we assume the Zeeman energy fluctuations are dominated by the Overhauser field rather than noise in the g factor difference.
The entangling two-qubit gate achieved between the spin qubits under the exchange interaction depends on the ratio $J/\Delta E_{z}$, tending to a $\sqrt{\text{SWAP}}$ operation as $J\gg\Delta E_{z}$ or a C-PHASE when $J\ll\Delta E_{z}$, though any gate within this set parameterised by $J/\Delta E_{z}$ can be used as the building block for a quantum error correcting code such as the surface code [63]. Defining the qubit quality factor as $Q=T_{2}^{*}\Omega$ (i.e. the number of periods before the amplitude of oscillations decays by $1/e$), we find $Q\gtrapprox 10$ in the region $J/\Delta E_{z}=2.1-3.2$, on par with prior reports across a range of semiconductor platforms [26, 47, 48, 58, 64, 65]. This provides an upper bound estimate on the achievable two-qubit gate fidelity using the approximation $\mathcal{F}\approx 1-1/4Q\lessapprox 98\%$ [66]. To implement error-correctable two-qubit gates this fidelity would need to surpass $99\%$ [15], which could be achieved using isotopically enriched silicon. In the next section, we extend the coherence time using spin refocusing techniques.
Fig. 5: |||| Echo sequence. a, Bloch sphere representation of the odd-parity two spin sub-space, indicating the rotation axes, $\hat{Z^{\prime}}$ and $\hat{X^{\prime}}$ and their angular deviation from the nominal $\hat{Z}$ axis defined by |$\ket{S}$ and |$\ket{T_{0}}$. b, Schematic of the exchange echo sequence. c, Echo signal as a function of free evolution time difference $\tau_{2}-\tau_{1}$. d, Echo amplitude as a function of total free evolution time $\tau_{2}=\tau_{1}$. Data (purple dots) with standard deviation and exponential decay fit (orange line).
@@
Fig. 6: |||| Extending radio-frequency cascade readout to two-dimensional arrays. a, Schematic representation of the cascade process extended to an arbitrarily long one-dimensional array. Pauli spin blockade readout is achieved by tunnel coupling the data QD (orange) and ancilla QA qubit (light purple), the resulting DQD is capacitively coupled to a chain of DQDs tuned into the cascade configuration (QCN) (dark purple). Enabling in-situ dispersive readout at an arbitrary distance. b, Two-dimensional array schematic unit-cell centred around a readout reservoir (yellow), connected to a radio-frequency reflectometry readout apparatus. Simultaneous multiplexed readout is achieved by applying distinct rf frequencies $f_{1}$ and $f_{2}$ to separate electrostatic gates, each controlling data qubit QDs, driving distinct cascaded charge transitions. c Repetition of the unit cell highlighting the robustness of the scheme to points of failure, such as a faulty readout reservoir, by re-routing the readout cascade chain to the nearest reservoir.
@@
In the exchange echo sequence, after initialising a |$\ket{\text{S}(1,1)}$ state and applying a $X^{\prime}_{\pi/2}$ rotation, the two-electron system dephases under the effect of charge noise for a time $\tau_{1}$. The free evolution occurs at a detuning point where $J/\Delta E_{z}=12.8\pm 1.6$ where we measured $T_{2}^{*}=43\pm 3~{}$ns (see Fig 4(f)). We then refocus the spins by applying a $X^{\prime}_{\pi}$ rotation and let the system evolve for $\tau_{2}$ until a second $X^{\prime}_{\pi/2}$ rotation maps the resulting state to the |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ axis. We extract the amplitude of the echo by fitting the signal in the $\tau_{2}-\tau_{1}$ domain (see Fig. 5(c)), and plot its value as a function of total free evolution time $\tau_{1}+\tau_{2}$. Fig. 5(d) shows the echo amplitude decays exponentially with total time ($\tau_{2}+\tau_{1}$), yielding a characteristic $T_{2,J}^{\text{echo}}=0.42\pm 0.02~{}\upmu$s which corresponds to an order of magnitude increase in the coherence time, similar to prior work with silicon devices [67, 26, 64]. From our fit to Eq. (4), we extract a magnetic-noise-limited $T_{2,\Delta E_{z}}^{*}=3.3~{}\upmu$s at this set point, indicating that sources other than magnetic noise limit $T_{2,J}^{\text{echo}}$. We speculate that residual high frequency charge noise limits the echo coherence [7], and may include factors such as the effect of the rf tone which is on throughout the control sequence.
@@ Discussion
In conclusion, we have demonstrated exchange control, which forms the basis for two-qubit gates between spin qubits, in an MOS DQD device fabricated using a 300 mm wafer process on natural silicon. The quality of the manufacturing manifests in the low detuning noise and relatively long $T_{2}^{*}$ for a natural silicon MOS device. These results highlight the advantages of industrial manufacturing and encourage follow-up studies in isotopically enriched Si samples, which, during the review of our manuscript, have now been conducted in Ref. [68]. Additionally, building on previous works [21], adapting dedicated gates to primarily control exchange strength over a wider range could (i) enable symmetric exchange pulses and (ii) reduce $J/\Delta E_{z}$ well below 1, which should lead to an overall reduction in sensitivity to charge noise.
Furthermore, we have expanded the portfolio of dispersive readout methods by introducing the radio-frequency electron cascade, which, with its enhanced gain, demonstrates highest performance for dispersive readout in planar MOS devices. Regarding scalability, the radio-frequency driven electron cascade can be extended to larger arrays for distant readout, eliminating the need for swap-based schemes that rely on shuttling information to a sensor [69]. As shown in Fig. 6, this extension builds upon existing schemes for two-dimensional grids using data and ancilla qubits [15, 4]. In this schematic, we assume that the tunnel barriers between each QD can be precisely controlled, enabling optimal readout fidelity through tuning $t_{c}$ [51]. Figure 6(a) illustrates the readout protocol in a simplified one-dimensional array. In order to readout the data qubit QD at the periphery of the unit cell, the data qubit is tunnel-coupled to a neighbouring ancilla qubit QA. The ancilla qubit, in turn, is capacitively coupled to a chain of DQDs configured in the cascade configuration, eventually connected to a reservoir. An rf-drive applied to the ancilla-data DQD initiates the cascade, propagating along the chain to a readout reservoir. This scheme uniquely enables the simultaneous readout of distant data qubits through frequency multiplexing at the readout tank circuit. In contrast to previous work [44], by driving multiple cascade chains at distinct frequencies, multiple distant qubits can be readout simultaneously, as shown in Fig. 6(b). Furthermore, the combination of unit cells into a dense, scalable 2D grid allows readout resource sharing, resulting in the system being resilient to failure points such as faulty QDs or reservoirs, as demonstrated in Fig. 6(c).
@@ Methods
[…]
Minimum integration time and readout fidelity calculation. The SNR of the interdot charge transition shown in Fig. 2(b) is determined by fitting the signal to an expression proportional to the quantum capacitance $\propto\partial^{2}E/\partial\varepsilon^{2}$ [42, 40],
$V_{\text{I,Q}}\propto V_{\text{I,Q}}^{0}-S_{\text{I,Q}}\frac{(2t_{c})^{3}}{\left((\varepsilon-\varepsilon_{0})^{2}+(2t_{c})^{2}\right)^{\frac{3}{2}}},$ (5)
where $t_{c}$ is the tunnel coupling, $\varepsilon_{0}$ is the detuning value of the centre of the peak, and $S_{\text{I,Q}}$ and $V_{\text{I,Q}}^{0}$ are the voltage signal and voltage offset in phase and quadrature components of the measured rf-response, respectively. The power SNR is determined by combining the measured in-phase an quadrature voltage signal components of the reflected radio-frequency signal,
$\text{SNR}=\frac{S_{I}^{2}+S_{Q}^{2}}{0.25(\sigma_{I}^{2}+\sigma_{Q}^{2})}$ (6)
where $S_{{\rm rf,I(Q)}}$ is indicated in Fig. 2(b), and $\sigma_{\rm rf,I(Q)}$ is the standard deviation in background signal (away from zero detuning). Here, we assume a white noise spectrum dominated by the first cryogenic amplification stage [70]. In the measurements acquired throughout the text, the applied rf-power is $P_{\text{rf}}=-78~{}$dBm, corresponding to a power $\text{SNR}=1010\pm 40$. The minimum integration time is determined using [40],
$\tau_{\text{min}}=\frac{N_{\text{avg}}}{2\Delta f~{}\text{SNR}}$ (7)
where $N_{\text{avg}}=4000$ is the number of averages used in the measurement, and $\Delta f=1.53f_{\text{LPF}}$ is the measurement bandwidth set by the low-pass filter cutoff frequency $f_{\text{LPF}}=100~{}$kHz. Together, these two parameters give the noise equivalent integration time $\tau_{\rm NE}=13~{}{\rm ms}$ for the measurement in Fig. 2(b). Substituting the SNR into eq. (7) we find a minimum integration time $\tau_{\text{min}}=13.0\pm 0.5~{}\upmu$s. In combination with the relaxation time $T_{1}$ extracted in Fig. 2(d), the readout fidelity in eq. (2) can be calculated.
Control sequence and charge noise estimation. The pulse sequence implemented to demonstrate exchange control is as follows:
  1. 1.
    E: Empty quantum dot $\text{Q}_{1}$ by biasing the gate voltages such that the ground state is in the (0,1) charge configuration, over a duration of 100 ns.
  2. 2.
    E-I: Initialise the ground state |$\ket{\text{S}}$ via an adiabatic ramp from the (0,1) to (0,2) configuration, over a duration of $10~{}\upmu$s.
  3. 3.
    I-S: Non-adiabatic pulse across the |$\ket{\text{S}}$-|$\ket{\text{T}_{-}}$ anti-crossing to $\varepsilon=0$, over a duration of 100100100~{}100ns. Establishing a symmetric detuning point to apply subsequent ramps.
  4. 4.
    S-P: Adiabatic ramp with respect to $\Delta E_{z}$ to $\varepsilon=0.926~{}$meV over a duration of 250 ns. Initialising $\ket{\downarrow\uparrow}$ where $J<\Delta E_{z}$.
  5. 5.
    P-J-P: Non-adiabatic pulse to near zero detuning $\varepsilon_{J}$ and back where $J\gg\Delta E_{z}$, for variable duration $\tau_{J}$.
  6. 6.
    P-M: Reverse adiabatic ramp to $\varepsilon=0$, projecting the final state onto |$\ket{\text{S}}$ or |$\ket{\text{T}_{0}}$ for readout.
  7. 7.
    M: Pre-measure delay of $6~{}\upmu$s is waited over before integrating for a measurement time of $8~{}\upmu$s.
The total duty cycle of the control sequence, $T_{\text{rep}}\approx 25~{}\upmu$s, provides the high frequency $f_{\text{high}}$ bound we integrate over to estimate $S_{0}$, the power spectral density noise at 111~{}1Hz. In combination with the total time to acquire a trace of exchange oscillations, $T_{M}=1/f_{\text{low}}=0.3~{}$h, we can estimate $S_{0}$ from [71],
[…]
where the power spectral density of charge noise has the functional form $S_{\varepsilon}(f)=S_{0}/f^{\alpha}$ with typical values of $\alpha$ ranging from $1$ to $2$. Considering the measured integrated charge noise ($\delta\varepsilon_{rms}=5.4~{}\upmu$eV), we estimate an upper and lower bound of $\sqrt{S_{0}^{\alpha=1}}=0.91~{}\upmu$eV$/\sqrt{\text{Hz}}$ and $\sqrt{S_{0}^{\alpha=2}}=0.12~{}\upmu$eV$/\sqrt{\text{Hz}}$ respectively. This result is on par with the state-of-the-art reported in planar MOS devices [36] and strained Ge wells (Ge/SiGe) [61] $\sqrt{S_{0}}=0.6\pm 0.3~{}\upmu\text{eV}/\sqrt{\rm Hz}$, as well as Si wells in Si/SiGe heterostructures $\sqrt{S_{0}}=0.3-0.8~{}\upmu{\rm eV}/\sqrt{\rm Hz}$ [64, 62].
Extended Data Fig. S1: |||| Tuning the radio-frequency driven electron cascade. a-e, Charge stability diagrams as a function of the DQD gate voltages $V_{\text{G}_{1}}-V_{\text{G}_{2}}$. The multi-electron QD bias $V_{G_{\text{S}}}$ is varied in each panel, defined with respect to $V_{\text{G}_{\text{S}}}^{\text{casc}}$, the bias where the conditions for cascade are observed, shown in panel c. This is the bias utilized in the main text of the manuscript. f-j, Simulated charge stability diagrams using the methods described in Supplementary Section S4. k-o Schematic depiction of the electrochemical levels for the corresponding bias conditions, arranged column-wise.
@@
S1 ELECTROSTATIC REQUIREMENTS FOR RF-DRIVEN ELECTRON CASCADE READOUT
The conditions for rf-driven electron cascade are similar to those stated in a previous report on electron cascade with a proximal charge sensor [44]. It can be understood by considering the electrochemical potentials $\mu_{i}$ for each quantum dot Qi, as well as the given charge configuration $(\text{N}_{\text{Q}_{1}},\text{N}_{\text{Q}_{2}},\text{N}_{\text{Q}_{\text{ME}}})$ where $\text{N}_{\text{Q}_{i}}$ refers to the number of charges in dot $\text{Q}_{i}$, as shown in Fig. 1 of the main text. For cascade to occur the DQD must be tuned to the $\text{Q}_{1}-\text{Q}_{2}$ inter-dot charge transition, such that
@@
Tuning the quantum dot array into rf-driven electron cascade requires precise control of the electrochemical potentials of the QDs, to the order of 10s of microvolts in the device presented here. The tuning procedure is most clearly demonstrated in the $V_{\text{Q}_{1}}-V_{\text{Q}_{2}}$ gate-voltage-space, as shown in Supplementary Fig. S1 where in each panel $V_{\text{G}_{\text{S}}}$ is varied, bringing the system into and out of cascade. In this configuration, the range of $V_{\text{G}_{\text{S}}}$ bias voltages satisfying the conditions for cascade are given by $V_{\text{G}_{\text{S}}}^{casc}=728.6\pm 0.1~{}$mV, a relatively narrow range of voltages as compared to the 141414~{}14mV addition voltage of $\text{Q}_{\text{ME}}$.
@@
In panels (f-j), we present matching radio-frequency simulations of the triple QD system that highlight the enhanced intensity of the interdot charge transition (see Section. S4). Further, we supplement the explanation with schematics of the electrochemical levels in each $V_{\text{G}_{\text{S}}}$ conditions in panels (k-o).
S2 AMPLIFICATION FACTOR
Here, we obtain the expression for the signal amplification factor generated by the cascade process, i.e. Eq. (1) in the main text. To determine the amplification factor, we consider two different charge movement events as seen from the electrode connected to the resonator, in this case, the electron reservoir (R):
@@
To obtain the amplification factor, we first consider the expressions for the signal-to-noise ratio for case (1) [40]. In particular, we assume that the system is in the low frequency limit $hf_{\text{rf}}\ll\Delta_{\text{c}}$ (where $f_{\text{rf}}$ is the frequency of the resonator and $\Delta_{\text{c}}=2t_{c}$), the small signal regime $Q_{\text{L}}\Delta C_{\text{Q}}/(2C_{\text{tot}})\ll 1$ (where $Q_{\text{L}}$ is the loaded quality factor of the resonator, $\Delta C_{\text{Q}}$ is the change in quantum capacitance of the system and $C_{\text{tot}}$ the total capacitance of the system) and the large excitation regime, $\alpha_{21}eV_{\text{dev}}\gg\Delta_{\text{c}}$ to ensure an electron tunnels every half cycle of the rf excitation. Here $\alpha_{21}=\alpha_{\text{R,2}}-\alpha_{\text{R,1}}$ is the interdot lever arm as seen from the reservoir and $V_{\text{dev}}$ is the amplitude of the oscillatory voltage arriving at the reservoir. In this case, the SNR is
@@
Here, the resonator is coupled to $Q_{\text{ME}}$ via a tunnel barrier, rather than a capacitively coupled gate, and hence the induced charge is $(1-\alpha_{\text{r,ME}})e$.
Supplementary Fig. S2: |||| Schematic circuit depicting alternating current induced by the radio-frequency driven electron cascade. Two tunneling processes are depicted between tunnel coupled (1) quantum dots Q1 and Q2 and (2) QME and the reservoir. The induced alternating current at the reservoir from each of these tunneling events is given by $I_{\rm R,i}=e\alpha_{R,i}f_{0}$ where $i=1,2,{\rm ME}$ and $\alpha_{\rm R,i}$ are the corresponding lever arms, with rf drive frequency $f_{0}$. The rf modulation applied to the system is represented by $V_{\rm dev}$, the device voltage.
@@
From this we can determine an expression for the amplification factor by considering the two tunneling processes presented in Supplementary Fig. S2. In process (1) the alternating current generated by cyclical tunneling from Q1 to Q2 is,
@@
In process (2), in addition to the current produced by process (1), we have the current produced by the cascade process adding to a total of,
@@
The amplification factor $A$ is then given by the ratio between the two processes,
@@
S3 MAGNETOSPECTROSCOPY SIMULATION
Here, we describe the simulations of the magnetospectroscopy map in the inset of Fig. 1(c) in the main text. that allow us to estimate both the tunnel coupling $t_{\text{c}}$ and the electron temperature $T_{\text{e}}$. We utilize the simplified Hamiltonian:
@@
Here, $T_{\text{e}}$ is the DQD temperature and $Z$ is the partition function over all states [72]. We plot the results of the simulations in Supplementary Fig. S3 where, in panel (a) and (b), we show the energy spectrum of the system as a function of detuning for $B=0.5$ T and 0 T, respectively, and in panel (c), we plot the normalised quantum capacitance of the systems as a function of $\varepsilon$ and $B$. We find that the best match between the data and simulations occurs when $t_{\text{c}}=2.4$ GHz and $T_{\text{e}}=50$ mK.
Supplementary Fig. S3: |||| Magneto-spectroscopy simulation of the (0,2)-(1,1) inter-dot charge transition. Eigenenergies of the two-electron spin states, singlet |$\ket{\text{S}}$ and triplet |$\ket{\text{T}}$ states as a function of detuning $\varepsilon$ normalised by tunnel coupling $\Delta=2t_{c}$. For a, applied magnetic field $B=0~{}$mT and b $B=500~{}$mT. c Normalised quantum capacitance simulation versus detuning and magnetic field, with the degeneracy point of the |$\ket{\text{S}}$-|$\ket{\text{T}_{-}}$ crossing overlaid (white dashed line).
S4 CHARGE STABILITY DIAGRAM SIMULATION
@@
$C_{cc}=\begin{pmatrix}35.000&-4.882&-1.886\\ -4.882&27.936&-0.402\\ -1.886&-0.402&42.048\end{pmatrix}$ (22)
@@
$C_{cv}=\begin{pmatrix}0.898&0.146&0.000\\ 0.537&0.453&0.000\\ 0.063&0.020&2.016\end{pmatrix}$ (23)
where each element is given in aF. The number of charges on each dot are given by $\vec{N}$, and $\vec{V_{G}}$ denotes the applied gate voltages. Throughout the simulations, we use natural units and set the charge of the electron $e=1$. We note that the ratio between capacitances in the matrices defined here are found to be in qualitative agreement with the data in Supplementary Fig. S1(a-e), however the magnitude of these capacitances have not been verified experimentally.